OpenStudy (anonymous):

What is the sum of the geometric sequence -4, 24, -144 .. if there are 8 items. I know they are going up by -12.

OpenStudy (anonymous):

no its -6.

OpenStudy (callisto):

common ratio = 24/-4 = -6 no. of terms = 8 first term = -4 \[Sum = \frac{a(1-r^n)}{1-r} = \frac{(-4)(1-(-6)^8)}{1-(-6)} = ...?\]